# Learning Admissible Heuristics for A*: Theory and Practice

### v1

@@ 1 Introduction
Heuristic search algorithms such as A* (Hart et al., 1968) are widely used in pathfinding tasks (Yonetani et al., 2021; Meng et al., 2024; Kirilenko et al., 2023), where the objective is to find a least-cost path from a start state to a goal state in a graph. These algorithms rely on a heuristic function that estimates the cost-to-go from a state to the goal. To guarantee optimality, A* requires the heuristic to be admissible, meaning it must never overestimate the true cost, $h^{*}$, to the goal. If we wish to integrate machine learning with heuristic search algorithms like A* we have several possible approaches: (1) We can adapt learning to meet the requirements of search algorithms (Samadi et al., 2008; Li et al., 2022), (2) we can adapt search algorithms to meet the properties of ML heuristics (Lelis et al., 2021), (3) we can do a combination of both, or (4) we can abandon all theoretical guarantees, hoping for high quality solutions or fast search in practice (Orseau et al., 2023; Hazra et al., 2024; Chen et al., 2025; Huang et al., 2025). This paper studies the foundations of the first question: what is the sample complexity needed to learn admissible heuristics, what is an effective loss function for training, and how close we can get to optimal heuristics in practice?
Traditionally, admissible heuristic functions are computed from domain knowledge of the problem (Katz & Keyder, 2022; Haslum et al., 2005; Seipp, 2024). One prominent example is pattern databases (PDBs) (Korf, 1997; Culberson & Schaeffer, 1998). PDBs result from abstracting the problem and storing $h^{*}$ for all states in the abstract problem. Learning heuristics from data (Li et al., 2022; Agostinelli et al., 2021; Pándy et al., 2022; Shen et al., 2020; Chen et al., 2024b) has become increasingly popular for two main reasons: (1) designing good heuristics for complex environments is challenging, and (2) data-driven deep learning algorithms have demonstrated outstanding performance in analogous tasks (Silver et al., 2017; Schrittwieser et al., 2020; Wurman et al., 2022; Touvron et al., 2023). While learned heuristic functions may outperform traditional heuristics, they typically lose admissibility and thus suboptimality guarantees (Agostinelli et al., 2019; Yonetani et al., 2021; Archetti et al., 2021; Kirilenko et al., 2023). Despite substantial empirical progress, the theoretical foundations of heuristic learning are relatively underexplored. A central question is: How many training samples are required to ensure that the learned heuristic generalizes effectively to the entire graph? Sakaue & Oki (2022) addressed this question for Greedy Best-First Search (GBFS) and A*, providing upper and lower bounds on the training samples necessary to ensure generalization.
@@ 2 Background
In heuristic search, the task is to find a path from a start state $s_{init}\in V$ to a goal state $goal\in V$ in a graph $\{G=\{V,E\},c,h\}$, where $c\colon E\rightarrow\mathbb{R}^{+}$ specifies the cost for each edge and the heuristic function $h(v)$ estimates the distance from any state $v$ to goalgoal. A heuristic is called admissible if, for all $v\in V$, we have $h(v)\leq h^{*}(v)$, where $h^{*}(v)$ is the true shortest distance from $v$ to goalgoal. It is consistent if, for any states $a,b\in V$, the inequality $h(a)\leq c(a,b)+h(b)$ holds. In large state spaces, the graph $G$ is represented implicitly and is generated dynamically during search by expanding states and exploring their neighbors. Although A* guarantees optimal solutions under an admissible heuristic, it can become less efficient if the heuristic is inconsistent, due to the potential for re-expanding nodes (Martelli, 1977; Felner et al., 2011; Helmert et al., 2019). For a positive integer $d$, write $d]\coloneqq\{1,2,\ldots,d\}$. Given vectors $\{\mathbf{x}^{1},\ldots,\mathbf{x}^{t}\}\subseteq\mathbb{R}^{d}$, we use superscripts to index the vectors and subscripts for coordinates, so $\mathbf{x}^{i}_{j}$ denotes coordinate $j$ of vector $i$. For any real vector $\mathbf{x}$, the operator $\lfloor\mathbf{x}\rfloor$ applies the floor function coordinatewise. We use $\mathbb{I}(\cdot)$ to denote the indicator function, which returns 1 when the stated condition holds and 0 otherwise. Let $\Pi$ denote the space of problem instances and let $D$ be an unknown distribution supported on $\Pi$; we write $x\sim D$ for a random instance. Each $x\in\Pi$ specifies a start state $s_{init}$ from which the search is initiated. We make the following assumption on all instances.
@@
As a form of abstraction heuristics, Pattern Databases (PDBs) simplify a graph $V$ by applying a homomorphic transformation to obtain a reduced state space $\phi(V)$. For example, a $3\times 3$ Rubik’s Cube contains 26 cubies—12 edges, 8 corners, and 6 centers—as illustrated in Figure 1. One can abstract away the edges and centers to construct an 8-corner PDB with $8!\times 3^{7}$ states, in contrast to the $4.33\times 10^{19}$ states in the original graph. Within this abstract representation, if there is an edge between $v_{1}$ and $v_{2}$ in $V$, an edge will also be present between $\phi(v_{1})$ and $\phi(v_{2})$ in $\phi(V)$. Thus, all distances are admissible, and all states from $V$ that reduce to the same state in $\phi(V)$ have identical heuristic values. More details regarding the reduction from states in $V$ to states in $\phi(V)$ are shown in Figure 2 in Appendix C. During the search, the current state is mapped to $\phi(v)$, and a ranking function (Myrvold & Ruskey, 2001) is used to assign a unique integer to the abstract state. This functions as an index in the lookup table to retrieve the the precomputed distance to the abstract goal, which serves as the heuristic value $h(v)$. Compression techniques (Felner et al., 2007; Helmert et al., 2017) reduce the size of the PDB by grouping entries and only storing the lowest heuristic value from the group, which maintains admissibility.
@@
We use pseudo-dimension (Pollard, 1984), a central concept for measuring the complexity of a class of real-valued functions, as the foundation of our analysis. The definition of pseudo-dimension is as follows.
@@
We use the following proposition (Theorem 11.8 in (Mohri et al., 2018)) that helps us to attain generalization bounds using the definition of pseudo-dimension.
@@
$\biggl|\frac{1}{N}\sum_{i=1}^{N}h(y_{i})\;-\;\operatorname{\mathbb{E}}_{y\sim\mathcal{D}}\bigl[h(y)\bigr]\biggr|=\mathrm{O}\!\biggl(H\sqrt{\frac{\mathrm{Pdim}(\mathcal{H})\,\lg\!\bigl(\tfrac{N}{\mathrm{Pdim}(\mathcal{H})}\bigr)+\lg\!\bigl(\tfrac{1}{\delta}\bigr)}{N}}\biggr).$
In simpler terms, if the size of the training dataset is $N=\Omega\left(\frac{H^{2}}{\epsilon^{2}}\left(\mathrm{Pdim}(\mathcal{H})\lg\frac{H}{\epsilon}+\lg\frac{1}{\delta}\right)\right)$, with a probability of $1-\delta$, we can guarantee that the difference between true expectation over the entire graph and the emperical mean over the dataset is less than $\epsilon$.
@@
Assumption 3 (Sakaue & Oki (2022, Assumption 3)).
@@
With this definition, the utility function could, for instance, be the number of node expansions or the total running time, since both can be directly computed from the sizes of the OPEN and CLOSED lists. As shown in Proposition 1, we require a strict upper bound on these measures to ensure that all functions in the class $\mathcal{H}$ return positive, bounded values. This assumption is explicitly used in experiments, as we often set time or memory constraints when running $A^{*}_{h}$. We denote by $u_{h}:\Pi\to[0,H]$ the utility function that evaluates $A^{*}_{h}$’s performance on any instance $x\in\Pi$, and we define the class of such functions as $\mathcal{U}=\{\,u_{h}:\Pi\to[0,H]\;|\;h\in\mathbb{R}^{n}\}.$
@@ 3 Sample Complexity Analysis
In this section, we attain the $\mathrm{Pdim}(\mathcal{U})$ for the $A^{*}_{h}$ algorithm. To this aim, we need to discretize $\mathbb{R}^{n}$ into regions such that all heuristics in a region have the same performance: $u_{h_{1}}=u_{h_{2}}$ if $h_{1},h_{2}\in\mathcal{P}$. If we look at the Algorithm 1, the only part that the heuristic function impacts the performance is the node selection step from OPEN (line 3). However, in addition to $h$, the path cost from $s_{init}$ to each state $v$ also matters to determine the total cost: $f(v)=g(v)+h(v)$. If only $h$ determines which node to expand, such as in GBFS, we know that only the total ranking given by $h$ over all states $v\in V$ matters. So, the number of regions are the total number of different ordering among the states, which is $n!$ (Sakaue & Oki, 2022; Chrestien et al., 2024). In the case of A* algorithm, we first need to define the cases when two different heuristics result in the same performance:
@@
Under Assumption 4, the number of distinct values that $g(v)$ can take for any $v\in V$ is at most $|V|=n$. Define $\mathcal{G}_{V}\;=\;\bigl\{\,(v,g(v))\mid v\in V\bigr\},$ which is the set of vertex–cost pairs. Hence, $|\mathcal{G}_{V}|=n^{2}$. Armed with these observations, we now establish an upper bound on $\mathrm{Pdim}(\mathcal{U})$ for A* in the following theorem.
@@
Sakaue & Oki (2022) provided the first upper bounds on $\mathrm{Pdim}(\mathcal{U})$. For a general class of graphs, they proved an upper bound of $\mathcal{O}(n^{2}\log n)$, later refining it to $\mathcal{O}(n\log(nW))$ under the assumption that edge weights are bounded by a constant $W$. In our setting, where graphs follow Assumption 4, we initially obtain an upper bound of $\mathcal{O}(n\log n)$. We then improve this bound to $\mathcal{O}(m\log n)$ by utilizing learned PDB heuristics, with $m$ representing the size of the graph induced by the PDB.
@@ 3.1 Upper bounds on the expected suboptimality
If we aim to achieve $\mathrm{Pdim}(\mathcal{U})$ for the A* algorithm without imposing Assumption 4, the sample complexity scales on the order of $\mathcal{O}(n^{2}\log n)$. Consequently, to obtain tighter results, one often restricts the class of graphs under consideration. Rather than seeking tighter bounds for a general class of performance measures, we focus on a special measure called the expected suboptimality, which captures how far the solution returned by A* is from the optimal solution. For a given problem instance $x\in\Pi$, let $C(x)$ be the cost of the path found by A*, and let $C^{*}(x)$ be the cost of an optimal path. We define the suboptimality as
@@
To reduce the error term in Equation (1), we employ a worst-case bound on the quality of the solution returned by A*. Valenzano et al. (2014) proved that if A* does not reopen nodes (i.e., if we remove lines 1314 in Algorithm 1), then the suboptimality can be bounded by the total inconsistency along the edges of an optimal path. Concretely, for any instance $x\in\Pi$, if $v_{0},v_{1},\dots,v_{k}$ is an optimal path with cost $C^{*}(x)$, then
@@
where, for every edge $(p,c)\in E$, the inconsistency $\mathrm{INC}_{h}$ is defined as $\mathrm{INC}_{h}(p,\,c)\;=\;\max\bigl\{\,h(p)\;-\;h(c)\;-\;c(p,c),\;0\bigr\}$. Because $\mathrm{INC}_{h}(\cdot,\cdot)$ is always non-negative, the total inconsistency on any path can be estimated by summing the edge inconsistencies along that path. Sakaue & Oki (2022) further showed that this bound remains valid even when A* reopen nodes from CLOSED upon discovering a path with lower $g$ cost. In Theorem 3, we show that the bound can be tightened when A* is allowed to reopen nodes.
@@
For the class $\hat{\mathcal{U}}$ of inadmissibility functions, it holds that $\mathrm{Pdim}(\hat{\mathcal{U}})=\mathrm{O}(n\lg n)$ and we can bound the generalization error by
$\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathcal{D}}\bigl[\Psi_{h}(x)\bigr]\;\;\leq\;\;\frac{1}{N}\sum_{i=1}^{N}\Psi_{h}(x_{i})\;+\;\tilde{\mathrm{O}}\Bigl(\hat{H}\sqrt{\tfrac{\,n+\lg\tfrac{1}{\delta}}{\,N}}\,\Bigr).$ (2)
@@
$\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathcal{D}}\left[C_{h}(x)\;-\;C^{*}(x)\right]\leq\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathcal{D}}[\Psi_{h}(x)]\leq\frac{1}{N}\sum_{i=1}^{N}\Psi_{h}(x_{i})+\tilde{\mathrm{O}}\left(\hat{H}\sqrt{\frac{n+\lg\frac{1}{\delta}}{N}}\right).$ (3)

### v2

@@ 1 Introduction
Heuristic search algorithms such as A* (hart1968formal) are widely used in pathfinding tasks (yonetani2021path; DBLP:conf/emnlp/Meng0YPC24; kirilenko2023transpath), where the objective is to find a least-cost path from a start state to a goal state in a graph. These algorithms rely on a heuristic function that estimates the cost-to-go from a state to the goal. To guarantee optimality, A* requires the heuristic to be admissible, meaning it must never overestimate the true cost, $h^{*}$, to the goal. If we wish to integrate machine learning with heuristic search algorithms like A* we have several possible approaches: (1) We can adapt learning to meet the requirements of search algorithms (samadi08; li2022optimal), (2) we can adapt search algorithms to meet the properties of ML heuristics (Lelis_Valenzano_Nazar_Stern_2021), (3) we can do a combination of both, or (4) we can abandon all theoretical guarantees, hoping for high quality solutions or fast search in practice (ijcai2023p624; hazra2024saycanpay; chen2025language; huang2025chasing). This paper studies the foundations of the first question: what is the sample complexity needed to learn admissible heuristics, what is an effective loss function for training, and how close we can get to optimal heuristics in practice?
Traditionally, admissible heuristic functions are computed from domain knowledge of the problem (katz2022search; haslum2005new; seipp2024dissecting). One prominent example is pattern databases (PDBs) (korf1997finding; culberson1998pattern). PDBs result from abstracting the problem and storing $h^{*}$ for all states in the abstract problem. Learning heuristics from data (li2022optimal; agostinelli2021obtaining; pandy2022learning; shen2020learning; chen2024return) has become increasingly popular for two main reasons: (1) designing good heuristics for complex environments is challenging, and (2) data-driven deep learning algorithms have demonstrated outstanding performance in analogous tasks (silver2017mastering; schrittwieser2020mastering; wurman2022outracing; touvron2023llama). While learned heuristic functions may outperform traditional heuristics, they typically lose admissibility and thus suboptimality guarantees (agostinelli2019solving; yonetani2021path; archetti2021neural; kirilenko2023transpath). Despite substantial empirical progress, the theoretical foundations of heuristic learning are relatively underexplored. A central question is: How many training samples are required to ensure that the learned heuristic generalizes effectively to the entire graph? sakaue2022sample addressed this question for Greedy Best-First Search (GBFS) and A*, providing upper and lower bounds on the training samples necessary to ensure generalization.
@@ 2 Background
In heuristic search, the task is to find a path from a start state $s_{init}\in V$ to a goal state $goal\in V$ in a graph $\{G=\{V,E\},c,h\}$, where $c\colon E\rightarrow\mathbb{R}^{+}$ specifies the cost for each edge and the heuristic function $h(v)$ estimates the distance from any state $v$ to goalgoal. A heuristic is called admissible if, for all $v\in V$, we have $h(v)\leq h^{*}(v)$, where $h^{*}(v)$ is the true shortest distance from $v$ to goalgoal. A heuristic function $h$ is consistent if, for all states $a,b\in V$ such that $(a,b)\in E$, we have $h(a)\leq c(a,b)+h(b)$. In large state spaces, the graph $G$ is represented implicitly and is generated dynamically during search by expanding states and exploring their neighbors. Although A* guarantees optimal solutions under an admissible heuristic, it can become less efficient if the heuristic is inconsistent, due to the potential for re-expanding nodes (martelli1977complexity; felner2011inconsistent; helmert2019ibex). For a positive integer $d$, write $d]\coloneqq\{1,2,\ldots,d\}$. Given vectors $\{\mathbf{x}^{1},\ldots,\mathbf{x}^{t}\}\subseteq\mathbb{R}^{d}$, we use superscripts to index the vectors and subscripts for coordinates, so $\mathbf{x}^{i}_{j}$ denotes coordinate $j$ of vector $i$. For any real vector $\mathbf{x}$, the operator $\lceil\mathbf{x}\rceil$ applies the ceil function coordinatewise. We use $\mathbb{I}(\cdot)$ to denote the indicator function, which returns 1 when the stated condition holds and 0 otherwise. Let $\Pi$ denote the space of problem instances and let $D$ be an unknown distribution supported on $\Pi$; we write $x\sim D$ for a random instance. Each $x\in\Pi$ specifies a start state $s_{init}$ from which the search is initiated. We make the following assumption on all instances.
@@
As a form of abstraction heuristics, Pattern Databases (PDBs) simplify a graph $V$ by applying a homomorphic transformation to obtain a reduced state space $\phi(V)$. For example, a $3\times 3$ Rubik’s Cube contains 26 cubies—12 edges, 8 corners, and 6 centers—as illustrated in Figure 1. One can abstract away the edges and centers to construct an 8-corner PDB with $8!\times 3^{7}$ states, in contrast to the $4.33\times 10^{19}$ states in the original graph. Within this abstract representation, if there is an edge between $v_{1}$ and $v_{2}$ in $V$, an edge will also be present between $\phi(v_{1})$ and $\phi(v_{2})$ in $\phi(V)$. Thus, all distances are admissible, and all states from $V$ that reduce to the same state in $\phi(V)$ have identical heuristic values. More details regarding the reduction from states in $V$ to states in $\phi(V)$ are shown in Figure 3 in Appendix C. The PDB stores the perfect heuristic value (optimal distance to the abstract goal) for every abstract state, obtained by running an optimal search in the abstract state space. During the search, the current state is mapped to $\phi(v)$, and a ranking function (myrvold2001ranking) is used to assign a unique integer to the abstract state. This functions as an index in the lookup table to retrieve the the precomputed distance to the abstract goal, which serves as the heuristic value $h(v)$. Compression techniques (felner2007compressed; helmert2017variable) reduce the size of the PDB by grouping entries and only storing the lowest heuristic value from the group, which maintains admissibility.
@@
We use pseudo-dimension (Pollard1984-zp), a central concept for measuring the complexity of a class of real-valued functions, as the foundation of our analysis. The definition of pseudo-dimension is as follows.
@@
We use the following proposition (Theorem 11.8 in (Mohri2018-zs)) that helps us to attain generalization bounds using the definition of pseudo-dimension.
@@
$\biggl|\frac{1}{N}\sum_{i=1}^{N}h(y_{i})\;-\;\operatorname{\mathbb{E}}_{y\sim\mathcal{D}}\bigl[h(y)\bigr]\biggr|=\mathrm{O}\!\biggl(H\sqrt{\frac{\mathrm{Pdim}(\mathcal{H})\,\log\!\bigl(\tfrac{N}{\mathrm{Pdim}(\mathcal{H})}\bigr)+\log\!\bigl(\tfrac{1}{\delta}\bigr)}{N}}\biggr).$
In simpler terms, if the size of the training dataset is $N=\Omega\left(\frac{H^{2}}{\epsilon^{2}}\left(\mathrm{Pdim}(\mathcal{H})\log\frac{H}{\epsilon}+\log\frac{1}{\delta}\right)\right)$, with a probability of $1-\delta$, we can guarantee that the difference between true expectation over the entire graph and the emperical mean over the dataset is less than $\epsilon$.
@@
Assumption 3 (sakaue2022sample).
@@
With this definition, the utility function could, for instance, be the number of node expansions or the total running time, since both can be directly computed from the sizes of the OPEN and CLOSED lists. As shown in Proposition 1, we require a strict upper bound on these measures to ensure that all functions in the class $\mathcal{U}$ return positive, bounded values. This assumption is explicitly used in experiments, as we often set time or memory constraints when running $A^{*}_{h}$. We denote by $u_{h}:\Pi\to[0,H]$ the utility function that evaluates $A^{*}_{h}$’s performance on any instance $x\in\Pi$, and we define the class of such functions as $\mathcal{U}=\{\,u_{h}:\Pi\to[0,H]\;|\;h\in\mathbb{R}^{n}\}.$
@@ 3 Sample Complexity Analysis
In this section, we attain the $\mathrm{Pdim}(\mathcal{U})$ for the $A^{*}_{h}$ algorithm. To this aim, we need to discretize $\mathbb{R}^{n}$ into regions such that all heuristics in a region have the same performance: $u_{h_{1}}=u_{h_{2}}$ if $h_{1},h_{2}\in\mathcal{P}$. If we look at the Algorithm 1, the only part that the heuristic function impacts the performance is the node selection step from OPEN (line 3). However, in addition to $h$, the path cost from $s_{init}$ to each state $v$ also matters to determine the total cost: $f(v)=g(v)+h(v)$. If only $h$ determines which node to expand, such as in GBFS, we know that only the total ranking given by $h$ over all states $v\in V$ matters. So, the number of regions are the total number of different ordering among the states, which is $n!$ (sakaue2022sample; chrestien2024optimize). In the case of A* algorithm, we first need to define the cases when two different heuristics result in the same performance:
@@
Under Assumption 4, the number of distinct values that $g(v)$ can take for any $v\in V$ is at most $|V|=n$. Define $\mathcal{G}_{V}\;=\;\bigl\{\,(v,g(v))\mid v\in V\bigr\},$ which is the set of vertex–cost pairs. Hence, $|\mathcal{G}_{V}|\leq n^{2}$. Armed with these observations, we now establish an upper bound on $\mathrm{Pdim}(\mathcal{U})$ for A* in the following theorem. While our proof uses a different derivation technique, the resulting bound matches that of sakaue2022sample.
@@
sakaue2022sample provided the first upper bounds on $\mathrm{Pdim}(\mathcal{U})$. For a general class of graphs, they proved an upper bound of $\mathcal{O}(n^{2}\log n)$, later refining it to $\mathcal{O}(n\log(nW))$ under the assumption that edge weights are bounded by a constant $W$. In our setting, where graphs follow Assumption 4, we initially obtain an upper bound of $\mathcal{O}(n\log n)$. We then improve this bound to $\mathcal{O}(m\log n)$ by utilizing learned PDB heuristics, with $m$ representing the size of the graph induced by the PDB.
@@ 3.1 Upper bounds on the expected suboptimality
If we aim to achieve $\mathrm{Pdim}(\mathcal{U})$ for the A* algorithm without imposing Assumption 4, the sample complexity scales on the order of $\mathcal{O}(n^{2}\log n)$. Consequently, to obtain tighter results, one often restricts the class of graphs under consideration. The general class of performance measures defined earlier encompasses a wide range of metrics related to the performance of the $A^{*}_{h}$ algorithm, including runtime, the number of node expansions, and memory consumption. Rather than seeking tighter bounds for this general class, we focus on a special measure called the expected suboptimality, which captures how far the solution returned by A* is from the optimal solution. For a given problem instance $x\in\Pi$, let $C(x)$ be the cost of the path found by A*, and let $C^{*}(x)$ be the cost of an optimal path. We define the suboptimality as
@@
To reduce the error term in Equation (1), we employ a worst-case bound on the quality of the solution returned by A*. valenzano2014worst proved that if A* does not reopen nodes (i.e., if we remove lines 1314 in Algorithm 1), then the suboptimality can be bounded by the total inconsistency along the edges of an optimal path. Importantly, this bound holds for any heuristic function—even if it is not admissible—as long as the following conditions are satisfied for all $s\in V$:
  • $h(s)\geq 0$,
  • $h(s)=0$ if $s=goal$, and
  • $h(s)\neq\infty$ whenever $h^{*}(s)\neq\infty$.
Concretely, for any instance $x\in\Pi$, if $v_{0},v_{1},\dots,v_{k}$ is an optimal path with cost $C^{*}(x)$, then
@@
where, for every edge $(p,c)\in E$, the inconsistency $\mathrm{INC}_{h}$ is defined as $\mathrm{INC}_{h}(p,\,c)\;=\;\max\bigl\{\,h(p)\;-\;h(c)\;-\;c(p,c),\;0\bigr\}$. Because $\mathrm{INC}_{h}(\cdot,\cdot)$ is always non-negative, the total inconsistency on any path can be estimated by summing the edge inconsistencies along that path. sakaue2022sample further showed that this bound remains valid even when A* reopen nodes from CLOSED upon discovering a path with lower $g$ cost. In Theorem 3, we show that the bound can be tightened when A* is allowed to reopen nodes.
@@
For the class $\hat{\mathcal{U}}$ of inadmissibility functions, it holds that $\mathrm{Pdim}(\hat{\mathcal{U}})=\mathrm{O}(n\log n)$ and we can bound the generalization error by
$\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathcal{D}}\bigl[\Psi_{h}(x)\bigr]\;\;\leq\;\;\frac{1}{N}\sum_{i=1}^{N}\Psi_{h}(x_{i})\;+\;\tilde{\mathrm{O}}\Bigl(\hat{H}\sqrt{\tfrac{\,n+\log\tfrac{1}{\delta}}{\,N}}\,\Bigr).$ (2)
@@
$\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathcal{D}}\left[C_{h}(x)\;-\;C^{*}(x)\right]\leq\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathcal{D}}[\Psi_{h}(x)]\leq\frac{1}{N}\sum_{i=1}^{N}\Psi_{h}(x_{i})+\tilde{\mathrm{O}}\left(\hat{H}\sqrt{\frac{n+\log\frac{1}{\delta}}{N}}\right).$ (3)